In a farm the number of chickens were 100% more than the cows. If there were 640 legs altogether how many chickens were there?
Answer by ikleyn(53763)

.

Let x be the number of cows.

Then the number of chickens is 2x, according to the condition.


Counting the legs, you have this equation


    4x + 2*(2x) = 640,   or


    4x + 4x = 640

      8x    = 640

       x    = 640/8 = 80.


ANSWER.  80 cows.

Solved, explained, answered and completed.

-------------

As a gift from me, consider these closely related lessons

    - Problem on two-wheel and three-wheel bicycles
    - Problem on animals at a farm
    - Problem on pills in containers
    - What type of problems are these?

in this site.